You and two friends of yours decided to handle the supply of vegetables to the shops of the city and created a limited liability company. Your friends have invested in the case 5 thousand euro each and you gave the company the right to use your personal car for one year. The car costs 4 thousand euro, you estimated your contribution in the form of the rights of use as agreed with the partners at 7 thousand euro. You have successfully completed the year of work. At your disposal 1800 euro of profit after accounting all costs. How will you share profit between partners?
1
Expert's answer
2019-11-11T15:24:43-0500

Share profit should be proportional to the contribution of capital participants. Total amount of all contribution is 5+5+7 = 17 thousands euro. In this way, both friends will receive 1800/17000*5000 = 529,41 euro each. And you will receive 1800/17000*7000 = 741.18 euro.